The perimeter of a rectangle is 86 cm the length is 19 cm greater than the width.
Answer by TimothyLamb(4379)

---
2x + 2y = 86
x = y + 19
---
put the system of linear equations into standard form:
---
2x + 2y = 86
x - y = 19
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 31 cm
y = 12 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php